# Do Language Models Have a Critical Period for Language Acquisition?

### v1

@@ Appendix E Derivation of EWC Regularization
where the transition from Eq. 3a to Eq. 3b follows from the conditional independence assumption of Eq. 2.
There is no known general algorithm to compute $p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}},\mathcal{D}_{\mathrm{L}_{2}})$. Thus, we construct a Gaussian approximation of $p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}})$ by computing a second-order Taylor approximation to the log-likelihood around the likelihood mode, i.e.,
@@
${\pi}(\boldsymbol{\theta})\propto\exp-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (6)
@@
$\log p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}})=\log p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta})+\log{\pi}(\boldsymbol{\theta})-\log p(\mathcal{D}_{\mathrm{L}_{1}})$ (7)
@@
$\approx\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\nabla^{2}_{\boldsymbol{\theta}}\log p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8b)
$=-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8c)
@@
$=-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(\sum_{n=1}^{N}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\boldsymbol{x}_{n}\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8d)
$=-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(N\cdot\left(\frac{1}{N}\sum_{n=1}^{N}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\boldsymbol{x}_{n}\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\right)\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8e)
@@
$\approx-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(N\cdot\operatorname*{\mathbb{E}}\left[\nabla^{2}_{\boldsymbol{\theta}}-\log p(\cdot\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\right]\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8f)
$=-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(N\cdot\mathbf{H}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8g)
A Fast Approximation to the Hessian.
@@
Computation of the matrix $\mathbf{H}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$ is impractical, although not intractable, for two reasons. First, it requires second-order automatic differentiation, which is relatively expensive. Second, it contains $\mathcal{O}(d^{2})$ unique entries. [^15] By Schwarz’s lemma, if a function is twice continuously differentiable, its Hessian is symmetric—hence, the big-O. When $d$ is large, as it is in our case, we cannot take its square. A classic identity involving the Fisher information matrix (Bickel and Doksum, 2001, pg. 185) suggests the following data-driven approximation to the Hessian:
@@
$\mathbf{H}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}\approx\mathbf{F}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}=\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim p(\cdot\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\mathbf{g}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}\mathbf{g}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}^{\intercal}$ (9)
where $\mathbf{g}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$ is in our case the gradient of the negative log-likelihood at $\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}$.
The quantity $\mathbf{F}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$ is known as the Fisher information matrix and is often obtained using an empirical estimation, where the expectation is calculated only based on the observed labels $\boldsymbol{x}_{n}$ from the dataset. However, Kunstner et al. (2019) have shown that such an estimation strategy lacks many useful properties of the real Fisher information, and in many cases can lead to inaccurate results. For this reason, we compute $\mathbf{F}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$ by sampling $K$ times [^16] The experimental choice of the number of samples $K$ is discussed in § 5 from our model distribution $p(\cdot\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$. Using the identity given in Eq. 9, and factoring in the $N$ from Eq. 8g, we calculate the Fisher matrix entries as follows: [^17] This means we calculate an unnormalized Fisher matrix, as we do not have to average over the log-likelihood values
@@
$\widetilde{\mathbf{F}}(\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})_{ii}=\frac{1}{NK}\sum_{n=1}^{N}\sum_{k=1}^{K}(\nabla_{\boldsymbol{\theta}}-\log p_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}(\boldsymbol{z}_{k}\mid\boldsymbol{x}_{n}))_{i}^{2}$ (10)
where $K$ is the number of samples $z$ taken from our $\mathrm{L}_{1}$ model distribution $p(\cdot\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$. Additionally, we have only approximated the diagonal of the Fisher matrix.
@@
[…]

### v2

@@ Appendix E Derivation of EWC Regularization for Language Modeling
where the transition from Eq. 3a to Eq. 3b follows from the conditional independence assumption of Eq. 2. There is no known general-purpose algorithm to compute $p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}},\mathcal{D}_{\mathrm{L}_{2}})$. Thus, we construct a Gaussian approximation of $p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}})$ by computing a second-order Taylor approximation to the log-likelihood around the likelihood mode, i.e.,
@@
${\pi}(\boldsymbol{\theta})\propto\exp-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta},$ (6)
@@
$\log p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}})=\log p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta})+\log{\pi}(\boldsymbol{\theta})-\log p(\mathcal{D}_{\mathrm{L}_{1}}).$ (7)
@@
$\log$ $p(\boldsymbol{\theta}\mid\mathcal{D}_{\mathrm{L}_{1}})\approx\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\nabla^{2}_{\boldsymbol{\theta}}\log p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8a)
@@
$=\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(\sum_{n=1}^{N}\left(\nabla^{2}_{\boldsymbol{\theta}}\log p(\textsc{eos}\mid\boldsymbol{x}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})+\sum_{t=1}^{T_{n}}\nabla^{2}_{\boldsymbol{\theta}}\log p(x_{t}^{(n)}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\right)\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta}$ (8b)
@@
$\approx-\frac{1}{2}(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}\left(\sum_{n=1}^{N}\sum_{t=1}^{T_{n}+1}\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\right)(\boldsymbol{\theta}-\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})-\frac{1}{2\sigma^{2}}\boldsymbol{\theta}^{\intercal}\boldsymbol{\theta},$ (8c)
where the last approximation replaces $\nabla^{2}_{\boldsymbol{\theta}}\log p(x_{t}^{(n)}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$ with its expectation $\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla^{2}_{\boldsymbol{\theta}}\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$ and performs a sign manipulation.
A Fast Approximation.
@@
Exact computation of the matrix $\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$ is impractical for multiple reasons. First, it requires second-order automatic differentiation, which is relatively expensive and not a standard feature in common automatic differentiation toolkits, e.g., PyTorch (Paszke et al., 2019). Second, recall that $p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$ is the next-symbol distribution of the language model, i.e., a distribution over $\overline{\Sigma}$, so $\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$ could be computed in $\mathcal{O}(|\overline{\Sigma}|)$ time. While linear, $\overline{\Sigma}$ is often large in practice in modern language models. Thirdly, the matrix contains $\mathcal{O}(d^{2})$ unique entries. [^20] By Schwarz’s lemma, if a function is twice continuously differentiable, its Hessian is symmetric—hence, the big-$\mathcal{O}$. When $d$ is large, as it is in our case, we cannot compute all $d^{2}$ entries easily. A classic identity involving the Fisher information matrix from Bickel and Doksum (2001, pg. 185) (see also Kunstner et al. (2019, Eq. 4)) allows us to derive the following approximation:
@@
$\sum_{n=1}^{N}\sum_{t=1}^{T_{n}+1}$ $\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla^{2}_{\boldsymbol{\theta}}-\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$
$\approx\sum_{n=1}^{N}\sum_{t=1}^{T_{n}+1}\operatorname*{\mathbb{E}}_{\overline{x}\sim p(\cdot\mid\boldsymbol{x}_{<t},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})}\nabla_{\boldsymbol{\theta}}\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\nabla_{\boldsymbol{\theta}}\log p(\overline{x}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}$ (9a)
@@
$\approx\sum_{n=1}^{N}\sum_{t=1}^{T_{n}+1}\frac{1}{K}\sum_{k=1}^{K}\nabla_{\boldsymbol{\theta}}\log p(\overline{x}^{(k)}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})\nabla_{\boldsymbol{\theta}}\log p(\overline{x}^{(k)}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})^{\intercal}$ (9b)
=$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\widetilde{\mathbf{F}}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$ (9c)
Eq. 9b is a standard Monte Carlo approximation where $\overline{x}^{(k)}\sim p(\cdot\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*})$. When $K\ll|\overline{\Sigma}|$, the sample-based approximation results in a significant speed-up. , to avoid $\mathcal{O}(d^{2})$ computation time, we only approximate the diagonal of $\widetilde{\mathbf{F}}_{\boldsymbol{\theta}_{\mathrm{L}_{1}}^{*}}$, which has $\mathcal{O}(d)$ entries.
@@
The second term corresponds exactly to the well-known $L_{2}$ regularization term resulting from the prior over the parameters $\boldsymbol{\theta}$. In practice, we generalize the coefficient $\frac{1}{2}$ into a tunable regularization coefficient $\lambda$ and the coefficient $\frac{1}{2\sigma^{2}}$ into a tunable regularization coefficient $\mu$. We tune $\lambda$ on held-out data, but set $\mu=0$ throughout the experiments and therefore omit it from the main text. [^21] Based on empirical observations this hyperparameter did not have an effect on the results.